# Comparing next-generation detector configurations for high-redshift gravitational wave sources with neural posterior estimation

### v1

@@
Abstract
The coming decade will be crucial for determining the final design and configuration of a global network of next-generation (XG) gravitational-wave (GW) detectors, including the Einstein Telescope (ET) and Cosmic Explorer (CE). In this study and for the first time, we assess the performance of various network configurations using neural posterior estimation (NPE) implemented in Dingo-IS–a method based on normalizing flows and importance sampling that enables fast and accurate inference. We focus on a specific science case involving short-duration, massive and high-redshift binary black hole (BBH) mergers with detector-frame chirp masses $\mathcal{M}_{\mathrm{d}}>100~\mathrm{M}_{\odot}$. These systems encompass early-Universe stellar and primordial black holes, as well as intermediate-mass black-hole binaries, for which XG observatories are expected to deliver major discoveries. Validation against standard Bayesian inference demonstrates that NPE robustly reproduces complex and disconnected posterior structures across all network configurations. For a network of two misaligned L-shaped ET detectors (2L MisA), the posterior distributions on luminosity distance can become multimodal and degenerate with the sky position, leading to less precise distance estimates compared to the triangular ET configuration. However, the number of sky-location multimodalities is substantially lower than the eight expected with the triangular ET, resulting in improved sky and volume localization. Adding CE to the network further reduces sky-position degeneracies, and the better performance of the 2L MisA configuration over the triangle remains evident.
@@ Abstract
show]filippo.santoliquido@gssi.it
I Introduction
@@
Next-generation (XG) detectors such as the Einstein Telescope (ET; Punturo:2010zz; ET:2019dnz; Branchesi:2023mws) and Cosmic Explorer (CE; Reitze:2019iox; Evans:2023euw; Gupta:2023lga) will shape the future of gravitational-wave (GW) astronomy. Two of the greatest scientific breakthroughs expected with XG detectors are the observation of early-Universe binary black hole (BBH) mergers, potentially detectable up to redshifts of $z\lesssim 100$, and the detection of intermediate-mass black hole (IMBH) binary mergers, with source-frame total masses reaching up to $\lesssim 10^{4}\,\mathrm{M}_{\odot}$ (Hall:2019xmm). High-redshift BBHs include two main types of systems: (i) BBHs formed from Population III (Pop. III) stars, which have been extensively and recently studied as promising high-redshift GW sources of astrophysical origin (Hartwig:2016nde; Belczynski:2016ieo; kinugawa2020; liubromm2020; tanikawa2022; wang2022; Mestichelli:2024djn; Santoliquido:2024oqs). In particular, in santoliquido2023 we found that between $\sim 20$% and $\sim 70$% of detectable Pop. III BBHs with ET could merge at $z>8$; (ii) Primordial black holes (PBHs), originating from the collapse of large inhomogeneities during the radiation era (Zeldovich:1967lct; Hawking:1974rv; Chapline:1975ojl; Carr:1975qj), are also expected to form at high redshift and span a wide mass spectrum, including the stellar-mass regime accessible to XG observatories, and a broad range of merger rates (Ivanov:1994pa; Guth:2012we; Ivanov:1997ia; Blinnikov:2016bxu; DeLuca:2020qqa; Franciolini:2022tfm; Ng:2022agi; Carr:2020gox; Franciolini:2023opt).
In the $10^{2}$$10^{3}$ M range, IMBHs are thought to form primarily through stellar collisions (2002ApJ...576..899P; 2015MNRAS.454.3150G; 2016MNRAS.459.3432M; 2021A&A...652A..54A; 2023MNRAS.526..429A), hierarchical black hole mergers (Gerosa:2017kvu; Fishbach:2017dwv; Gerosa:2021mno; 2002MNRAS.330..232C; 2004ApJ...616..221G; 2019MNRAS.486.5008A; 2021A&A...652A..54A; Kritos:2022ggc), close interactions between stars and black holes (2017MNRAS.467.4180S; 2023MNRAS.521.2930R; 2023MNRAS.526..429A), or a combination of these mechanisms (2015MNRAS.454.3150G; 2023MNRAS.526..429A).
@@
Accurate parameter estimation is essential for uncovering the origins of early-Universe BBHs and for determining the relative importance of the various formation pathways of IMBHs. However, the large expected detection rate of XG observatories, about $10^{5}$ events per year (Baibhav:2019gxm), poses unprecedented challenges for parameter estimation and data analysis (see also chapter 10 of Abac:2025saz).
@@
To handle the vast amount of data expected, it is crucial to accelerate parameter estimation for individual sources, as current stochastic sampling methods, often requiring several hours per event (Smith:2019ucc), are inadequate for the era of XG detectors. Acceleration of parameter estimation can be pursued through three complementary strategies: simplifying the likelihood function (2010arXiv1007.4820C; Veitch:2014wba; Zackay:2018qdy; 2021PhRvD.104l3030L; Narola:2023men; Vinciguerra:2017ngf; Aubin:2020goo; 2021PhRvD.104d4062M; Allene:2025saz; 2015PhRvL.114g1104C; Smith:2021bqc; 2023PhRvD.107h4037T; Morras:2023pug; Smith:2021bqc; Narola:2024qdh; 2024arXiv241202651H), improving sampling efficiency (Williams:2021qyt; 2023MLS&T...4c5011W; nessai; 2025MNRAS.541..200P; 2023JOSS....8.5021W; 2023ApJ...958..129W; 2024PhRvD.110h3033W; Perret:2025wox; Negri:2025cyc; Prathaban:2025qgg), and employing methods based on deep neural networks (Cuoco:2020ogp).
In recent years, likelihood-free inference (or simulation-based inference)—particularly neural posterior estimation (NPE, Green:2020hst; Green:2020dnx; Dax:2021myb; Chatterjee:2022ggk; 2023mla..confE..34W; 2023PhRvL.130q1402L; Dax:2022pxd; DeSanti:2024oap; Lanchares:2025kpb; Marx:2025ioo; Srinivasan:2025etu; Kofler:2025dux)—has enabled much faster and accurate parameter estimation, producing posterior samples within minutes.
@@
Dax:2021tsq implemented NPE in Dingo using conditional normalizing flows (Papamakarios:2019fms; Kobyzev_2021), which are invertible neural networks that transform a base distribution, typically a standard normal, into a more complex distribution which serves as an approximation of the posterior. Santoliquido:2025lot trained Dingo on the triangular configuration of ET and showed that NPE performs effectively in the XG-detector regime, delivering accurate parameter estimation with excellent computational efficiency.
@@
The science goals of XG observatories critically depend on design choices that are effectively irreversible, including the detectors’ geometry, geographic location, and orientation (e.g., Branchesi:2023mws; Gupta:2023lga). Making the right decisions at the design stage is therefore crucial to maximizing the scientific potential of XG detectors.
@@
This work presents the first quantitative assessment of how different XG detector configurations affect the parameter estimation performance for massive BBH mergers ($\mathcal{M}_{\mathrm{d}}>100~\mathrm{M}_{\odot}$), based on full Bayesian analyses. Leveraging NPE enables accurate evaluation of key metrics, including localization volumes and the presence of multimodal sky posteriors.
@@
This manuscript is structured as follows. Section II.1 outlines the examined detector configurations, while Section II.2 introduces Dingo-IS, which combines NPE with importance sampling. Section II.3 defines the performance metrics used throughout the work. Section III.1 shows that Dingo-IS correctly recovers a large fraction of injections, between 85% and 96%, depending on the detector configuration. Section III.2 presents a single-injection study showing that a standard inference method and Dingo-IS yield statistically indistinguishable posteriors, with the latter requiring far less computational time. Section III.3 compares parameter-estimation performance across detector configurations, highlighting that two misaligned L-shaped ET detectors outperform the triangular ET for sky localization and volume. Finally, Section IV summarizes our conclusions.
II Methods
II.1 Detector and network configurations
We consider seven different XG detector configurations:
@@ II.1 Detector and network configurations
  • $\Delta$. One single triangular ET with 10-km arms located in Sardinia (Branchesi:2023mws).
@@
  • 2L A. Two L-shaped ET detectors, each with 15-km arms, positioned in Sardinia and in the Meuse-Rhine Euro region (EMR), with parallel arms (Branchesi:2023mws).
@@
  • 2L MisA. Same as 2L A, but with the detector arms misaligned by 45° (Branchesi:2023mws).
@@
  • 2L MisA + LHI. Same as the 2L MisA configuration, but including three LIGO observatories—Livingston, Hanford, and India (Unnikrishnan:2013qwa)—operating at the mid-2030s A# LIGO sensitivity.
@@
  • 1L + CE. An L-shaped ET detector with 15-km arms in Sardinia combined with a single CE detector with 40-km arms located near Hanford, WA (Reitze:2019iox). This configuration, introduced in Maggiore:2024cwf but not adopted as an official design by the ET Collaboration, is included here solely for comparison purposes.
@@
  • $\Delta$ + CE. A single triangular ET detector located in Sardinia with 10-km arms observing with CE, as defined in 1L + CE.
@@
  • 2L MisA + CE. Same as 1L + CE, but with a second L-shaped ET detector located at EMR, rotated by 45° relative to the first ET.
@@
During the development of this work, the Lusatia region in Germany, near Kamenz (51.275°, 14.100°), was announced as the third official candidate site for ET. Its chord distance from the Sardinian site is approximately 1248 km, only about 82 km longer than the distance between the EMR region and Sardinia. Given this minimal difference, we consider the two sites effectively equivalent in hosting the second L-shaped interferometer, for the specific science case considered here.
@@
Additional details, including coordinates, chord distances between interferometers and further information on the amplitude spectral densities (ASDs) are provided in Appendix A.
II.2 Dingo-IS
@@
We conduct our analysis using NPE implemented in Dingo, where a conditional probabilistic neural network $q_{\phi}(\theta|d)$ with tunable parameters $\phi$ approximates the true posterior $p(\theta|d)$. Dingo is trained on simulated datasets $(\theta,d)$, where the parameters $\theta$ are sampled from the priors and the corresponding data $d$ consist of a GW signal embedded in stationary Gaussian noise, $d=h(\theta)+n$.
The prior distribution is the same as that adopted in Santoliquido:2025lot and listed in Appendix B, along with the choices of waveform parameters and approximants.
@@
Compared to the network architecture in Santoliquido:2025lot, where all Dingo hyperparameters were kept at their default values, we enlarge the embedding network that compresses the input strain data into a set of features. Specifically, the output dimensionality is increased from 128 to 256 features, and several additional hidden layers are added, resulting in a total of $4\times 10^{8}$ learnable parameters.
Appendix C shows the learning curves for each considered detector configuration, along with the corresponding probability–probability plots.
@@
After training, Dingo can rapidly generate approximate posterior samples. These samples, however, may deviate from the true posterior. A practical way to assess and refine them is via importance sampling (https://doi.org/10.1002/wics.56; Payne:2019wmy; Elvira_2021; Ashton:2025xba). Combining this technique with Dingo leads to Dingo-IS (Dax:2022pxd), which assigns weights to a set of $n$ samples $\theta_{i}$ drawn from the proposal distribution $q_{\phi}(\theta|d)$:
@@
where $\mathcal{L}(d|\theta)$ is the likelihood function defined in Appendix D. Ideally, all weights $w_{i}$ would be equal; in practice, the sample efficiency quantifies how well $q_{\phi}(\theta|d)$ approximates the target distribution:
@@
with the effective number of samples given by $n_{\mathrm{eff}}=(\sum_{i}w_{i})^{2}/\sum_{i}w_{i}^{2}$ (Kong). The lower bound $\epsilon=1/n$ corresponds to the extreme case where one weight dominates and all others are negligible, which is the minimum possible as the weights are normalized to have mean one.
We compare the results obtained with Dingo-IS to those derived using standard inference methods implemented in Bilby (see Appendix D for additional details).
II.3 Metrics to assess detector performance
We define the metrics to evaluate the performance in estimating parameters of the various configurations of XG detectors.
The information gain ($I$) is defined as the Kullback–Leibler divergence (KL; 10.1214/aoms/1177729694) between the prior and the posterior (2022RNAAS...6...89B). In particular, we compute the KL divergence of the posterior from the prior:

### v2

@@
11institutetext: Gran Sasso Science Institute (GSSI), I-67100 L’Aquila, Italy 22institutetext: INFN, Laboratori Nazionali del Gran Sasso, I-67100 Assergi, Italy 33institutetext: Marietta Blau Institute (MBI) - Austrian Academy of Sciences, $10$ Vienna, Austria
Comparing next-generation detector configurations for high-redshift gravitational wave sources
with neural posterior estimation
F. Santoliquido [Uncaptioned image], Corresponding author: filippo.santoliquido@gssi.it    J. Tissino [Uncaptioned image]    U. Dupletsa [Uncaptioned image]    M. Branchesi [Uncaptioned image]    J. Harms [Uncaptioned image]
The coming decade will be crucial for determining the final design and configuration of a global network of next-generation (XG) gravitational-wave detectors, including the Einstein Telescope (ET) and Cosmic Explorer (CE). In this study, and for the first time, we assessed the performance of various network configurations using neural posterior estimation (NPE) implemented in Dingo-IS—a method based on normalizing flows and importance sampling that enables fast and accurate inference. We focused on a specific science case involving short-duration, massive and high-redshift binary black hole mergers with detector-frame chirp masses $(\mathcal{M}_{\mathrm{d}})$ $100\penalty 10000\ \mathrm{M}_{\odot}$. These systems encompass early-Universe stellar and primordial black holes, as well as intermediate-mass black hole binaries, for which XG observatories are expected to deliver major discoveries. Validation against standard Bayesian inference demonstrates that NPE robustly reproduces complex and disconnected posterior structures across all network configurations. For a network of two misaligned L-shaped ET detectors (2L MisA), the posterior distributions on luminosity distance can become multimodal and degenerate with the sky position, leading to less precise distance estimates compared to the triangular ET configuration. However, the number of sky-location multimodalities is substantially lower than the eight expected with the triangular ET, resulting in improved sky and volume localization. Adding CE to the network further reduces sky-position degeneracies, and the better performance of the 2L MisA configuration over the triangle remains evident.
Key Words.:
gravitational waves – methods: statistical – stars: black holes
@@
1 Introduction
@@
Next-generation (XG) detectors such as the Einstein Telescope (ET; Punturo and others, 2010; Maggiore et al., 2020; Branchesi and others, 2023) and Cosmic Explorer (CE; Reitze and others, 2019; Evans and others, 2023; Gupta and others, 2024) will shape the future of gravitational-wave (GW) astronomy. Two of the greatest scientific breakthroughs expected with XG detectors are the observation of early-Universe binary black hole (BBH) mergers, potentially detectable up to redshifts $\lesssim 100$, and the detection of intermediate-mass black hole (IMBH) binary mergers, with source-frame total masses reaching up to $\lesssim 10^{4}\,\mathrm{M}_{\odot}$ (Hall and Evans, 2019). High-redshift BBHs include two main types of systems: (i) BBHs formed from Population III (Pop. III) stars have been extensively and recently studied as promising high-redshift GW sources of astrophysical origin (Hartwig et al., 2016; Belczynski et al., 2017; Kinugawa et al., 2020; Liu and Bromm, 2020; Tanikawa et al., 2022; Wang et al., 2022; Mestichelli et al., 2024; Santoliquido et al., 2024). In particular, in Santoliquido et al. (2023) we found that between $\sim 20$% and $\sim 70$% of Pop. III BBHs detectable with ET could merge at $z>8$. (ii) Primordial black holes (PBHs), which originate from the collapse of large inhomogeneities during the radiation era (Zel’dovich and Novikov, 1967; Hawking, 1974; Chapline, 1975; Carr, 1975), are also expected to form at high redshifts and span a wide mass spectrum, including the stellar-mass regime accessible to XG observatories, and a broad range of merger rates (Ivanov et al., 1994; Guth and Sfakianakis, 2012; Ivanov, 1998; Blinnikov et al., 2016; De Luca et al., 2020; Franciolini et al., 2022; Ng et al., 2022; Carr et al., 2021; Franciolini et al., 2023).
@@
In the $10^{2}$$10^{3}$ M range, IMBHs are thought to form primarily through stellar collisions (Portegies Zwart and McMillan, 2002; Giersz et al., 2015; Mapelli, 2016; Arca Sedda et al., 2021, 2023), hierarchical black hole mergers (Gerosa and Berti, 2017; Fishbach et al., 2017; Gerosa and Fishbach, 2021; Miller and Hamilton, 2002; Gültekin et al., 2004; Antonini et al., 2019; Arca Sedda et al., 2021; Kritos et al., 2024), close interactions between stars and black holes (Stone et al., 2017; Rizzuto et al., 2023; Arca Sedda et al., 2023), or a combination of these mechanisms (Giersz et al., 2015; Arca Sedda et al., 2023). Accurate parameter estimation is essential for uncovering the origins of early-Universe BBHs and for determining the relative importance of the various formation pathways of IMBHs. However, the high expected detection rate of XG observatories, about $10^{5}$ events per year (Baibhav et al., 2019), poses unprecedented challenges for parameter estimation and data analysis (see also chapter 10 of Abac and others, 2025c).
@@
To handle the vast amount of data expected, it is crucial to accelerate parameter estimation for individual sources, as current stochastic sampling methods, often requiring several hours per event (Smith et al., 2020), are inadequate for the era of XG detectors. The acceleration of parameter estimation can be pursued through three complementary strategies: simplifying the likelihood function (Cornish, 2010; Veitch and others, 2015; Zackay et al., 2018; Leslie et al., 2021; Narola et al., 2024; Vinciguerra et al., 2017; Aubin and others, 2021; Morisaki, 2021; Alléné and others, 2025; Canizares et al., 2015; Smith and others, 2021; Tissino et al., 2023; Morras et al., 2023; Smith and others, 2021; Narola and others, 2025; Hu and Veitch, 2025), improving sampling efficiency (Williams et al., 2021, 2023; Williams, 2021; Prathaban et al., 2025a; Wong et al., 2023a, b; Wouters et al., 2024; Perret et al., 2025; Negri and Samajdar, 2025; Prathaban et al., 2025b), and employing methods based on deep neural networks (Cuoco and others, 2021).
@@
In recent years, likelihood-free inference (or simulation-based inference)—particularly neural posterior estimation (NPE; Green et al., 2020b; Green and Gair, 2021; Dax et al., 2021a; Chatterjee et al., 2023; Wildberger et al., 2023; Langendorff et al., 2023; Dax et al., 2023; De Santi et al., 2024; Lanchares et al., 2025; Marx et al., 2025; Srinivasan et al., 2025; Kofler et al., 2025; Chan et al., 2026)—has enabled much faster and accurate parameter estimation, producing posterior samples within minutes. Dax et al. (2021b) implemented NPE in Dingo using conditional normalizing flows (Papamakarios et al., 2021; Kobyzev et al., 2021), which are invertible neural networks that transform a base distribution, typically a standard normal, into a more complex distribution, which serves as an approximation of the posterior. Santoliquido and others (2025) trained Dingo on the triangular configuration of ET and showed that NPE performs effectively in the XG-detector regime, delivering accurate parameter estimation with excellent computational efficiency.
@@
The science goals of XG observatories critically depend on design choices that are effectively irreversible, including the detectors’ geometry, geographic location, and orientation (e.g., Branchesi and others, 2023; Gupta and others, 2024). Making the right decisions at the design stage is therefore crucial to maximizing the scientific potential of XG detectors.
@@
This work presents the first quantitative assessment of how different XG detector configurations affect the parameter estimation performance for massive BBH mergers ($\mathcal{M}_{\mathrm{d}}>100\penalty 10000\ \mathrm{M}_{\odot}$), based on full Bayesian analyses. Leveraging NPE enables the accurate evaluation of key metrics, including localization volumes and the presence of multimodal sky posteriors.
@@
This manuscript is structured as follows. Section 2.1 outlines the examined detector configurations, while Sect. 2.2 introduces Dingo-IS, which combines NPE with importance sampling. Section 2.3 defines the performance metrics used throughout the work. Section 3.1 shows that Dingo-IS correctly recovers a large fraction of injections, between 85% and 96% depending on the detector configuration. Section 3.2 presents a single-injection study showing that a standard inference method and Dingo-IS yield statistically indistinguishable posteriors, with the latter requiring far less computational time. Section 3.3 compares parameter-estimation performance across detector configurations, highlighting that two misaligned L-shaped ET detectors outperform the triangular ET for sky localization and volume. Finally, Sect. 4 summarizes our conclusions.
2 Methods
2.1 Detector and network configurations
We considered seven different XG detector configurations:
@@ 2.1 Detector and network configurations
  • $\Delta$. One single triangular ET with 10 km arms located in Sardinia (Branchesi and others, 2023).
@@
  • 2L A. Two L-shaped ET detectors, each with 15 km arms, positioned in Sardinia and in the Meuse-Rhine Euro (EMR) region, with parallel arms (Branchesi and others, 2023).
@@
  • 2L MisA. Same as 2L A but with the detector arms misaligned by 45° (Branchesi and others, 2023).
@@
  • 2L MisA + LHI. Same as the 2L MisA configuration but including three LIGO observatories—Livingston, Hanford, and India (Unnikrishnan, 2013)—operating at the mid-2030s A# LIGO sensitivity.
@@
  • 1L + CE. An L-shaped ET detector with 15 km arms in Sardinia combined with a single CE detector with 40 km arms located near Hanford (Reitze and others, 2019). This configuration, introduced in Maggiore et al. (2025) but not adopted as an official design by the ET Collaboration, is included here solely for comparison purposes.
@@
  • $\Delta$ + CE. A single triangular ET detector located in Sardinia with 10 km arms observing with CE, as defined in 1L + CE.
@@
  • 2L MisA + CE. Same as 1L + CE but with a second L-shaped ET detector located in the EMR region, rotated by 45° relative to the first ET.
@@
During the development of this work, the Lusatia region in Germany, near Kamenz (51.275°, 14.100°), was announced as the third official candidate site for ET. Its chord distance from the Sardinian site is approximately 1248 km, only about 82 km longer than the distance between the EMR region and Sardinia. Given this minimal difference, we consider the two sites effectively equivalent for hosting the second L-shaped interferometer for the specific science case considered here.
@@
Additional details, including coordinates, chord distances between interferometers, and further information on the amplitude spectral densities (ASDs), are provided in Appendix A.
2.2 Dingo-IS
@@
We conducted our analysis using NPE implemented in Dingo, where a conditional probabilistic neural network ($q_{\phi}(\theta|d)$) with tunable parameters ($\phi$) approximates the true posterior ($p(\theta|d)$). Dingo is trained on simulated datasets $(\theta,d)$, where the parameters ($\theta$) are sampled from the priors and the corresponding data ($d$) consist of a GW signal embedded in stationary Gaussian noise, $d=h(\theta)+n$. The prior distribution is the same as that adopted in Santoliquido and others (2025) and is listed in Appendix B, along with the choices of waveform parameters and approximants.
@@
Compared to the network architecture in Santoliquido and others (2025), where all Dingo hyperparameters were kept at their default values, we enlarged the embedding network that compresses the input strain data into a set of features. Specifically, the output dimensionality was increased from 128 to 256 features, and several additional hidden layers were added, resulting in a total of $4\times 10^{8}$ learnable parameters. Appendix C shows the learning curves for each considered detector configuration, along with the corresponding probability–probability plots.
@@
After training, Dingo can rapidly generate approximate posterior samples. These samples, however, can deviate from the true posterior. A practical way to assess and refine them is via importance sampling (Tokdar and Kass, 2010; Payne et al., 2019; Elvira and Martino, 2021; Ashton, 2025). Combining this technique with Dingo leads to Dingo-IS (Dax et al., 2023), which assigns weights to a set of $n$ samples of $\theta_{i}$ drawn from the proposal distribution $q_{\phi}(\theta|d)$:
@@
where $\mathcal{L}(d|\theta)$ is the likelihood function defined in Appendix D. Ideally, all weights ($w_{i}$) would be equal; in practice, the sample efficiency quantifies how well $q_{\phi}(\theta|d)$ approximates the target distribution:
@@
with the effective number of samples given by $n_{\mathrm{eff}}=(\sum_{i}w_{i})^{2}/\sum_{i}w_{i}^{2}$ (Kong, 1992). The lower bound $\epsilon=1/n$ corresponds to the extreme case where one weight dominates and all others are negligible, which is the minimum possible as the weights are normalized to have a mean equal to 1. We compared the results obtained with Dingo-IS to those derived using standard inference methods implemented in Bilby (see Appendix D for additional details).
2.3 Metrics to assess detector performance
We defined the metrics to evaluate the performance in estimating parameters of the various configurations of XG detectors. The information gain ($I$) is defined as the Kullback–Leibler (KL) divergence (Kullback and Leibler, 1951) between the prior and the posterior (Buchner, 2022). In particular, we computed the KL divergence of the posterior from the prior: